A truck traveling at a constant rate of 45 miles per hour leaves Albany. One hour later a car traveling at a constant rate of 60 miles per hour also leaves Albany traveling in the same direction of the same highway. How long will it take for the car to catch up to the truck if both vehicles continue in the same direction on the highway? I do not understand can u please help?
Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12844)

D=RT
D=45T FOR THE TRUCK.
D=60(T-1) FOR THE CAR.
BECAUSE THE DISTANCES ARE EQUAL:
45T=60(T-1)
45T=60T-60
45T-60T=-60
-15T=-60
T=-60/-15
T=4 HOURS AFTER THE TRUCK LEAVES THE CAR WILL CATCH THE TRUCK.
PROOF:
45*4=60(4-1)
180=60*3
180=180

Answer by Alan3354(69443)

A truck traveling at a constant rate of 45 miles per hour leaves Albany. One hour later a car traveling at a constant rate of 60 miles per hour also leaves Albany traveling in the same direction of the same highway. How long will it take for the car to catch up to the truck if both vehicles continue in the same direction on the highway?
-----------
In 1 hour, the 1st truck has gone 45 miles.
The 2nd truck gains on it at 15 mph (60 - 45)
45/15 = 3 hours
------------
That's 3 hours after the car leaves, 4 hours after the truck has left.